A thermometer contains 3.00 g of mercury. After it gains 8.4 J of energy  the mercury has reached the 5.00°C mark. What was the initial  temperature of the mercury?


1
Expert's answer
2021-06-03T18:28:34-0400

Solution.

"m=3.00g;"

"Q=8.4J;"

"t=5.00^o;"

"c=130J\/kg^oC;"

"Q=cm(t-t_0);"

"cmt_0=cmt-Q\implies t_0=\dfrac{cmt-Q}{cm};"

"t_0=\dfrac{130J/kg^oC\sdot3.00\sdot10^{-3}kg\sdot5.00^oC-8.4J}{130J/kg^oC\sdot3.00\sdot10^{-3}kg}=-16.5^oC;"

Answer: "t_0=-16.5^oC."